# Parametric Mean-Field Empirical Bayes in High-dimensional Linear Regression

### v1

@@ 1 Introduction
While traditional Bayesian inference assumes that the prior $\mu_{\theta}$ is specified by the statistician, this can be problematic for high-dimensional noisy data. For example, assume the simple high-dimensional setting where $p=n$ and $X$ is the $n\times n$ identity matrix. This is the classical Gaussian sequence model
$y_{i}\overset{i.i.d.}{\sim}N(\beta_{i},\sigma^{2}).$
In this case, the posterior distribution of $\beta\mid y,X$, witnesses proportional contribution from the prior and the likelihood, which implies that there is no posterior contraction. In other words, all posterior based inference depends heavily on the prior, even asymptotically as $n\to\infty$. A misspecified prior from the statistician would therefore result in inconsistent results downstream. This motivates the development of procedures that can consistently estimate $\mu_{\theta_{0}}$. Empirical Bayes (EB) addresses this problem by using the observed data to estimate the prior. For simplicity, suppose the variance $\sigma^{2}$ is known. A natural strategy for estimating $\mu_{\theta_{0}}$ is to maximize the marginalized likelihood
@@
with respect to the ‘hyperparameter’ $\theta$. However, except in extreme cases (such as when $\mu_{\theta}$ is a normal distribution or $X^{\top}X$ is diagonal), the integral in (2) does not have an explicit form. The problem is particularly challenging under the high-dimensional setting with $p\to\infty$, as the marginal likelihood in (2) is extremely challenging to compute by standard numerical integration or sampling-based approaches.
Recent works (see [10, 27, 38, 20]) have addressed this issue by maximizing a more tractable variational lower bound, also commonly known as the evidence lower bound (ELBO), instead of (2), to obtain a variational empirical Bayes (vEB) estimator. Necessary conditions for the consistency of this estimator was recently established in [38]. However, to our knowledge, the explicit convergence rates, asymptotic distributions, and asymptotic efficiency of vEB estimators have not been studied in the literature. In this work, we take a first step towards this direction by analyzing an approximate vEB estimator in the parametric setting (see [10, 21, 22]). We provide a complete characterization of its asymptotic behavior when $p=o(n)$, under appropriate scaling of the design matrix. We then use the estimated prior for coordinate-wise and delocalized inference under the oracle posterior.
@@ 1.1 Organization and summary of main results
This work first analyzes frequentist inferential guarantees of the vEB estimator $\hat{\theta}^{\mathsf{vEB}}$ that maximizes a variational approximation of the marginal likelihood $m_{\theta}(y)$, assuming a random design matrix $X$ and regularity conditions for the prior parametric family. Interestingly, the optimality of $\hat{\theta}^{\mathsf{vEB}}$ exhibits a phase transition depending on the asymptotic scaling of the problem. When $p\ll n^{2/3}$, the vEB estimator is $\sqrt{p}$-consistent, asymptotically normal, and attains the information-theoretically optimal variance. In contrast, all of these properties fail when $p\gg n^{2/3}$. This is formally stated in the following theorem, which combines the separate upper and lower bound claims in Section 3. While the following result is stated under a correctly specified parametric family for the prior, we also derive analogous limit distributions under misspecified priors (see 3.4).
@@ 1.2 Related works
Standard EB in denoising.
The standard EB setup for normal means [this corresponds to taking $X=I_{n}$ in (1), see e.g., 46, 24, 25, 51] lead to a product posterior, and hence results in independent observations under the posterior. This allows the marginal likelihood to factorize, and does not require a high-dimensional integral computation as in (2). Additionally, the exact factorization of the posterior results in tractable representations for individual coordinates of the posterior mean and variance, using Tweedie’s formula (see e.g. [46]). In contrast, our regression setup does not have any such representations of the posterior moments due to the intractable likelihood, creating additional theoretical challenges for downstream inference.
Statistical efficiency of MML for EB.
As stated earlier in (2), maximizing the marginal likelihood is a natural approach to empirical Bayes. Consequently, asymptotic properties of parametric MML was previously studied in the literature [42, 53, 47, 45]. In particular, [42] established consistency of the MML estimator to an “oracle set” as well as the consistency of the EB posterior. The recent work [45] additionally shows a higher-order merging of the EB posterior to the oracle posterior, which indicates that the EB posterior credible intervals enjoy coverage under the oracle posterior. However, this line of work considered the more classical finite-dimensional asymptotic scaling with high SNR. More importantly, they posit a frequentist true model that corresponds to fixing any true parameter $\beta^{\star}$ in the regression case, thus making the Bayesian model formulation to be misspecified. This is in sharp contrast to our high-dimensional low SNR scaling with correct model specification. Indeed, we will show that such higher-order merging may not happen under our setup. We also stress that going beyond consistency, no asymptotic normality or efficiency results have previously been established for the MML estimator, beyond simple settings.
Gradient flow based methods.
@@
vEB for general latent variable models.
@@ 2 Variational empirical Bayes
For our goal of estimating $\theta$, the natural strategy is to maximize the marginal likelihood $m_{\theta}(y)$ (see (2)) with respect to $\theta$. We first define some transformations of the data $(X,y)$, which will be useful in simplifying $m_{\theta}(y)$.
@@
Using these notation, we expand the quadratic term in the exponent of (2):
@@
We motivate the vEB (variational empirical Bayes) estimator, initially assuming a known variance $\sigma^{2}>0$. While maximizing $\ell_{p}(\theta)$ as a function of $\theta$ is tempting, the normalizing constant $Z_{p}(w,\theta)$ in (5) involves an intractable high-dimensional integral which makes the maximization computationally challenging. To remedy this issue, we build our estimator of $\theta_{0}$ using ideas from Mean-Field variational inference. To wit, first observe that the Gibbs variational formula implies the following dual representation of $\log Z_{p}(w,\theta)$ (c.f. [59, 39]):
@@
The supremum in (8) is still an implicit function of $\theta$. We therefore approximate it further to reduce to a more tractable optimization problem in $\theta$.
@@
Under 3.1, by [34, Lemma 2.1], the optimizer attaining the supremum in (8) is unique. Let us denote it by $u_{\theta}=(u_{1,\theta},\ldots,u_{p,\theta})$. Then $u_{\theta}$ satisfies ${u}_{i,\theta}=\psi_{i,\theta}^{\prime}({s}_{i,\theta}+w_{i})$, where $s_{i,\theta}:=\sum_{j\neq i}A_{p}(i,j){u}_{i,\theta}$, for $i\leq p$.
For heuristics, assume that $\mu_{\theta}$ is compactly supported. Then under our random design condition (see 3.1 below), direct computations show that the empirical measure $p^{-1}\sum_{i=1}^{p}\delta_{s_{i,\theta}}$ converges weakly to 0. It is then natural to define the approximate optimizer $\tilde{u}_{\theta}=(\tilde{u}_{1,\theta},\ldots,\tilde{u}_{n,\theta})$ as $\tilde{u}_{i,\theta}:=\psi^{\prime}_{i,\theta}(w_{i}).$ As a result, by spelling out the KL divergences in $I_{i,\theta}$, we can write
@@
$F_{i}(\theta):=\log\int e^{w_{i}\beta_{i}-\frac{d_{i}\beta_{i}^{2}}{2}}d\mu_{\theta}(\beta_{i}).$ (9)
Note that $F_{i}(\theta)=C_{\theta}(d_{i})+\psi_{i,\theta}(w_{i})$. By plugging in the above observations into the lower bound (8), we have
$\ell_{p}(\theta)\geq\ell_{p}^{\text{MF}}(\theta)\approx-\frac{n}{2p}\log(2\pi\sigma^{2})-\frac{1}{2p\sigma^{2}}\lVert y\rVert^{2}+\frac{1}{p}F_{i}(\theta).$ (10)
@@ 3 Prior recovery via vEB and Fisher optimality
Next, we impose the following regularity conditions on the prior, which are standard assumptions for efficient estimation in parametric models [e.g. see Theorems 2.6 or 5.1 in 35]. For example, exponential families satisfy these conditions.
@@
Suppose that the prior $\mu_{\theta}$ is a parametric family with log-likelihood $\ell(\theta;\beta)$ with respect to a non-degenerate base measure $\nu$, such that satisfies the following regularity conditions:
@@ 3.1 Upper bound
    For $t\in\mathbb{R},~d>0$, let $B_{t,d,\theta}$ be a linear-quadratic tilted probability measure on $\mathbb{R}$ defined as $\frac{dB_{t,d,\theta}}{dB_{\theta}}(b)\propto e^{tb-\frac{db^{2}}{2}},$ so that $B_{\theta}\mid(W_{\theta}=t)\sim B_{t,d_{0},\theta}$. Finally, set $h_{\theta}:\mathbb{R}\to\mathbb{R}^{k}$ as $h_{\theta}(t):=\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{t,d_{0},\theta})$ and define a mapping $\kappa:\Theta\to\mathbb{R}^{k}$ as
    $\kappa(\theta):=\frac{1}{2}\operatorname{\mathbb{E}}[B_{\theta}^{2}]\operatorname{\mathbb{E}}[h_{\theta}^{\prime\prime}(W_{\theta})].$
@@
Let $\tilde{\ell}(\theta;\cdot)$ be the log-likelihood of the random variable $W_{\theta}$. Then, the corresponding Fisher information satisfies:
@@
    $\sup_{d\in B_{\epsilon}(d_{0})}\sup_{\theta\in\Theta}\big|\operatorname{\mathbb{E}}[B_{0,d,\theta}]\big|\lesssim 1,\quad\sup_{d\in B_{\epsilon}(d_{0})}\sup_{\theta\in\Theta}\|B_{t,d,\theta}\|_{\psi_{2}}\lesssim 1+|t|,$
@@
R4 is a technical polynomial growth condition for the first three derivatives of the log-likelihood. Note that this can be easily satisfied since this is a requirement for functions which are already in the log-scale. This is required to control the remainder term in the Taylor expansion for analyzing the first-order condition of (11), and is reminiscent of the usual integrability assumption for $\nabla^{3}\ell(\theta;\beta)$. Here we need the additional regularity for $\nabla\ell(\theta;\beta)$ and $\nabla^{2}\ell(\theta;\beta)$ as the third derivative of $F_{i}(\theta)$ involves all three derivatives. Also, R5 is a standard moment assumption for the prior. Note that it is weaker than a sub-Gaussian assumption on the prior.
@@
R6 requires sub-Gaussian tails for all linear and quadratic tilts of $B_{\theta}$. This is weaker than assuming sub-Gaussianity directly on the prior $\mu_{\theta}$. Here, as $\Theta$ and $B_{\epsilon}(d_{0})$ are bounded sets, the uniform control over all $\theta,d$ is natural. On the other hand, the linear tilt parameter $t$ is unbounded, so instead of taking a supremum over $t$, we allow the sub-Gaussian norm to depend on $|t|$. This flexibility allows us to cover log-concave as well as non log-concave mixture priors (see Lemma 3.2 below). We note that our results go through with minor modifications if we replace the linear growth of the sub-Gaussian norm in R6 with a general polynomial order growth condition. The linear growth is primarily for expositional simplicity.
The following lemma illustrates concrete instances where conditions R5, R6 in 3.4 hold, which already generalizes existing assumptions in the literature. In particular, we can accommodate priors with unbounded support, mixture priors, and sparsity inducing priors. On the other hand, existing works on low SNR high-dimensional Bayesian regression require the prior to have (a) bounded support [38, 33, 20], (b) log-concave densities [32, 49], or (c) densities that satisfy the log-Sobolev inequality alongside a stronger version of our R4 [21]. These assumptions in the existing literature exclude common choices in practical EB methods, such as the adaptive shrinkage priors (this is essentially scale mixtures of Gaussians, see [52]).
@@
    $\mu_{\theta}$ is a finite mixture of $\delta_{0}$ and Lebesgue densities which satisfy (R5), the first condition in (R6), and the stronger second condition
@@
    In particular, these conditions for the densities are satisfied by distributions from parts (a) and (b) above.
@@
Part (b) considers higher dimensions with $p\propto n^{2/3}$. While the convergence rate and limiting variance is same as that in part (b), an additional bias appears in the limit. Note that this bias grows as $\delta=\lim_{p\to\infty}\frac{p^{3/2}}{n}$ increases. In part (c), we further increase the dimension so that $\delta\to\infty$, or equivalently $n^{2/3}\ll p\ll n$. The resulting limiting distribution exhibits a phase transition and results in a slower rate of convergence (since $\frac{n}{p}\ll\sqrt{p}$ under this regime). This is because now the bias term in part (b) dominates the variance.
To better understand this phase transition, we briefly discuss the proof strategy when $k=1$. Assuming that $\theta$ is an one-dimensional parameter, the first-order conditions for (11) gives
@@
$\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X,\beta^{\star}\approx N\big(\mu_{p}(X,\beta^{\star}),T_{2}\big),\quad\text{where}\quad\mu_{p}(X,\beta^{\star}):=\frac{1}{\sqrt{p}}\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X,\beta^{\star}\Big]$ (16)
and $T_{2}>0$ is some constant variance (see (51)). The phase transition on $p$ arises from analyzing this mean term, where we use second-order Taylor expansions to write
$\mu_{p}(X,\beta^{\star})\approx\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\Phi(d_{0}\beta_{i}^{\star})+\frac{p^{3/2}}{n}\kappa(\theta_{0})$ (17)
for some function $\Phi$ with $\operatorname{\mathbb{E}}\Phi(d_{0}\beta_{i}^{\star})=0$ (see (52), (53)). When $p\ll n^{2/3}$, the first summand in (17) dominates and gives the limit $\mu_{p}(X,\beta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,T_{1})$. Combining this with the conditional normal limit (16) using the fact that $T_{1}+T_{2}=V(\theta_{0})$, we get $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,V(\theta_{0}))$. The conclusion in part (a) follows by Slutsky’s theorem.
@@
The convergence rate $\sqrt{p}$ is information-theoretically tight. To see this, consider the setting where one directly observes all regression coefficients $\beta^{\star}=(\beta_{1}^{\star},\ldots,\beta_{p}^{\star})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu_{\theta_{0}}$. The resulting convergence rate for estimating $\theta_{0}$ is $\sqrt{p}$, and hence a faster convergence rate is impossible when one is only given $y$.
@@
By summing up the separate limiting distributions of $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\hat{\beta}_{i}-\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}{\beta}^{\star}_{i}\mid y,X$, $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}{\beta}^{\star}_{i}-M(\theta_{0})\mid X$ and applying the Delta method, we can show the following limit:
@@
While this limit generally holds under $p=o(n)$, the limiting variance is sub-optimal and strictly greater than $V(\theta_{0})^{-1}$.
@@
In part (c) of 3.3, note that the $k$-dimensional vector $\kappa(\theta_{0})$ may be zero. While this singularity is typically a measure-zero event (see Figure 1 below for a visual illustration), the convergence rate at such points is faster than the $\frac{n}{p}$-rate in part (c). In particular, assuming $\kappa(\theta_{0})=0$, we can extend the proof of 3.3 to show that $\hat{\theta}^{\mathsf{vEB}}-\theta_{0}=O_{P}\Big(\frac{1}{\sqrt{p}}+\big(\frac{p}{n}\big)^{3/2}\Big)$. Consequently, (15) extends to the higher-dimensional asymptotic regime of $p=o(n^{3/4})$, and the phase transition no longer happens at $p\asymp n^{2/3}$.
@@
To illustrate the limits in 3.3, we consider a toy example of the Gaussian prior with unknown mean. Here, the conjugacy allows directly computing the limit distribution of $\hat{\theta}^{\mathsf{vEB}}$.
@@
As a toy example, suppose that the prior is a Gaussian with unknown mean and known variance: $\mu_{\theta}\stackrel{{\scriptstyle d}}{{\equiv}}N(\theta,1)$. Then, the integral for defining the functions $F_{i}$ becomes tractable (see (11)), and we can write
@@
By combining the two displays above and noting $\frac{1_{p}^{\top}\Sigma 1_{p}}{p}=d_{0}+O_{P,X}\Big(\frac{1}{\sqrt{n}}\Big)$ (standard mean and variance computations using the design matrix randomness), we have obtained the limit distribution of $\hat{\theta}^{\mathsf{vEB}}$:

### v2

@@ 1 Introduction
While traditional Bayesian inference assumes that the prior $\mu_{\theta}$ is specified by the statistician, this can be problematic for high-dimensional noisy data. For example, consider a toy scenario where $p=n$ and $X$ is the $n\times n$ identity matrix. This is the classical Gaussian sequence model
$y_{i}\mid\beta_{i}\overset{ind}{\sim}N(\beta_{i},\sigma^{2}).$
In this case, the posterior distribution of $\beta\mid y,X$ witnesses proportional contribution from the prior and the likelihood, which implies that there is no posterior contraction. In other words, all posterior based inference depends heavily on the prior, even asymptotically as $n\to\infty$. A misspecified prior from the statistician would therefore result in inconsistent posterior inference. This motivates the development of procedures that can consistently estimate $\mu_{\theta_{0}}$.
Empirical Bayes (EB) addresses this problem by using the observed data to estimate the prior. For simplicity, suppose the variance $\sigma^{2}$ is known. A natural strategy for estimating $\mu_{\theta_{0}}$ is to maximize the marginalized likelihood
@@
with respect to the ‘hyperparameter’ $\theta$. However, except in extreme cases (such as when $\mu_{\theta}$ is a normal distribution or $X^{\top}X$ is diagonal), the integral in (2) does not have an explicit form. The problem is particularly challenging under the high-dimensional setting with $p\to\infty$, as the marginal likelihood in (2) is extremely difficult to compute by standard numerical integration or sampling-based approaches.
Recent works (see [10, 29, 40, 21]) have addressed this issue by maximizing a more tractable variational lower bound, also commonly known as the evidence lower bound (ELBO), instead of (2). Commonly known as variational empirical Bayes (vEB), this procedure has been extremely popular in modern machine learning problems (such as Latent Dirichlet Allocation in [9] and variational autoencoders in [30]), but with limited theoretical understanding. While necessary conditions for consistency was recently established in [40], the explicit convergence rates, limiting distributions, and asymptotic efficiency of vEB estimators have not been studied in the literature, to our knowledge. In this work, we take a first step in this direction by analyzing an approximate vEB estimator in the parametric setting (see [10, 22, 23]). We provide a complete characterization of its asymptotic behavior when $p=o(n)$, under appropriate scaling of the design matrix. We then use the estimated prior for coordinate-wise and delocalized posterior inference.
@@ 1.1 Organization and summary of main results
This work first analyzes frequentist inferential guarantees of the vEB estimator $\hat{\theta}^{\mathsf{vEB}}$ that maximizes a variational approximation of the marginal likelihood $m_{\theta}(y)$, assuming a random design matrix $X$ and regularity conditions for the prior parametric family. Interestingly, the optimality of $\hat{\theta}^{\mathsf{vEB}}$ exhibits a phase transition depending on the asymptotic scaling of the problem. When $p\ll n^{2/3}$, the vEB estimator is $\sqrt{p}$-consistent, asymptotically normal, and attains the information-theoretically optimal variance. In contrast, all of these properties fail when $p\gg n^{2/3}$. This is stated in the following theorem, which combines the separate upper and lower bound claims in Section 3. While the following result is stated under a correctly specified parametric family for the prior, we also derive analogous limit distributions under misspecified priors (see 3.4).
@@ 1.2 Related works
1.2.1 Standard EB in denoising
The standard EB setup for normal means [this corresponds to taking $X=I_{n}$ in (1), see e.g., 48, 25, 26, 11, 54] lead to a product posterior, and hence results in independent observations under the posterior. This allows the marginal likelihood to factorize, and does not require a high-dimensional integral computation as in (2). Additionally, the exact factorization of the posterior results in tractable representations for individual coordinates of the posterior mean and variance, using Tweedie’s formula (see e.g. [48]). In contrast, our regression setup does not have any such representations of the posterior moments due to the intractable normalizing constant, creating additional theoretical challenges for downstream inference.
1.2.2 Statistical efficiency of MML for EB
As stated earlier in (2), maximizing the marginal likelihood (MML) is a natural approach to empirical Bayes. Consequently, asymptotic properties of parametric MML was previously studied in the literature [44, 56, 49, 11, 47, 66]. In particular, [44] established consistency of the MML estimator to an “oracle set” as well as the consistency of the EB posterior. The recent work [47] additionally shows a higher-order merging of the EB posterior to the oracle posterior, which indicates that the EB posterior credible intervals enjoy coverage under the oracle posterior. However, this line of work considers a more classical scaling with high SNR where the EB posterior contracts. This is in sharp contrast to our high-dimensional low SNR scaling. Indeed, we will show that such higher-order merging may not happen under our setup. We also stress that going beyond consistency, no asymptotic normality or efficiency results have previously been established for the MML estimator, beyond simple settings.
1.2.3 Gradient flow based methods
@@
1.2.4 vEB for general latent variable models
@@ 2 Variational empirical Bayes
To estimate the prior parameter $\theta$, we maximize a variational lower bound of the marginal log-likelihood $\log m_{\theta}(y)$. We first define some transformations of the data $(X,y)$, which will be useful in simplifying this quantity.
@@
Using these notations, we expand the quadratic term in the exponent of (2):
@@
We motivate the vEB (variational empirical Bayes) estimator, initially assuming a known variance $\sigma^{2}>0$ (see Remark 2.1 for the relaxation). While maximizing $\ell_{p}(\theta)$ as a function of $\theta$ is tempting, the normalizing constant $Z_{p}(w,\theta)$ in (5) involves an intractable high-dimensional integral which makes the maximization computationally challenging. To remedy this issue, we build our estimator of $\theta_{0}$ using ideas from Mean-Field variational inference. To wit, first observe that the Gibbs variational formula implies the following dual representation of $\log Z_{p}(w,\theta)$ (c.f. [62, 41]):
@@
The supremum in (8) is still an implicit function of $\theta$. We therefore approximate it further to reduce to a more tractable objective in $\theta$.
@@
Under a random design matrix (see 3.1 below), by [36, Lemma 2.1], the supremum in (8) has a unique optimizer. Let us denote it by $u_{\theta}=(u_{1,\theta},\ldots,u_{p,\theta})$. Then $u_{\theta}$ satisfies
${u}_{i,\theta}=\psi_{i,\theta}^{\prime}({s}_{i,\theta}+w_{i}),\quad\text{where}\quad s_{i,\theta}:=\sum_{j\neq i}A_{p}(i,j){u}_{j,\theta},\quad\forall i\leq p.$ (9)
Here, the identity (9) follows by writing the first order conditions of the supremum.
For heuristics, assume that $\mu_{\theta}$ is compactly supported. Then under our random design condition (see 3.1), direct computations show that the empirical measure $p^{-1}\sum_{i=1}^{p}\delta_{s_{i,\theta}}$ converges weakly to 0. It is then natural to approximate $s_{i,\theta}\approx 0$ in the identity ${u}_{i,\theta}=\psi_{i,\theta}^{\prime}({s}_{i,\theta}+w_{i})$ and define the approximate optimizer $\tilde{u}_{\theta}=(\tilde{u}_{1,\theta},\ldots,\tilde{u}_{n,\theta})$ as $\tilde{u}_{i,\theta}:=\psi^{\prime}_{i,\theta}(w_{i}).$ As a result, by spelling out the KL divergences in $I_{i,\theta}$, we can write
@@
$F_{i}(\theta):=\log\int e^{w_{i}\beta_{i}-\frac{d_{i}\beta_{i}^{2}}{2}}d\mu_{\theta}(\beta_{i})=C_{\theta}(d_{i})+\psi_{i,\theta}(w_{i}).$ (10)
By plugging in the above observations into the lower bound (8), we have
$\ell_{p}(\theta)\geq\ell_{p}^{\text{MF}}(\theta)\approx-\frac{n}{2p}\log(2\pi\sigma^{2})-\frac{1}{2p\sigma^{2}}\lVert y\rVert^{2}+\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}(\theta).$ (11)
@@ 3 Prior recovery via vEB and Fisher optimality
Next, we impose the following regularity conditions on the prior, which are standard assumptions for efficient estimation in parametric models [e.g. see Theorems 2.6 or 5.1 in 37]. For example, exponential families satisfy these conditions.
@@
Suppose that the prior $\mu_{\theta}$ is a parametric family with log-likelihood $\ell(\theta;\beta)$ with respect to a non-degenerate base measure $\nu$, such that it satisfies the following regularity conditions:
@@ 3.1 Upper bound
    For $t\in\mathbb{R},~d>0$, let $B_{t,d,\theta}$ be a linear-quadratic tilted probability measure on $\mathbb{R}$ defined as $\frac{dB_{t,d,\theta}}{dB_{\theta}}(b)\propto e^{tb-\frac{db^{2}}{2}},$ so that $B_{\theta}\mid(W_{\theta}=t)\sim B_{t,d_{0},\theta}$. Finally, set $G_{\theta}:\mathbb{R}\to\mathbb{R}^{k}$ as $G_{\theta}(t):=\operatorname{\mathbb{E}}\nabla\ell(\theta;B_{t,d_{0},\theta})$ and define a mapping $\kappa:\Theta\to\mathbb{R}^{k}$ as
    $\kappa(\theta):=\frac{1}{2}\operatorname{\mathbb{E}}[B_{\theta}^{2}]\operatorname{\mathbb{E}}[G_{\theta}^{\prime\prime}(W_{\theta})].$
@@
Let $\tilde{\ell}(\theta;\cdot)$ be the log-likelihood of the random variable $W_{\theta}$. Then, the corresponding Fisher information is $V(\theta)$:
@@
    $\sup_{d\in B_{\epsilon}(d_{0})}\sup_{\theta\in\Theta}\big|\operatorname{\mathbb{E}}[B_{0,d,\theta}]\big|\lesssim 1,\quad\sup_{d\in B_{\epsilon}(d_{0})}\sup_{\theta\in\Theta}\|B_{t,d,\theta}\|_{\psi_{2}}\lesssim 1+|t|,\quad\forall t\in\mathbb{R},$
@@
R4 is a technical polynomial growth condition for the first three derivatives of the log-likelihood. Note that this condition is mild since the log-likelihood is typically polynomially bounded in $\beta$ for common parametric families. This is required to control the remainder term in the Taylor expansion for analyzing the first-order condition of (12), and is reminiscent of the usual integrability assumption for $\nabla^{3}\ell(\theta;\beta)$. Here we need the additional regularity for $\nabla\ell(\theta;\beta)$ and $\nabla^{2}\ell(\theta;\beta)$ as the third derivative of $F_{i}(\theta)$ involves all three derivatives. Also, R5 is a standard moment assumption for the prior. Note that it is weaker than a sub-Gaussian assumption on the prior.
@@
R6 requires sub-Gaussian tails for linear and quadratic tilts of $B_{\theta}$. This is weaker than assuming sub-Gaussianity directly on the prior $\mu_{\theta}$. Here, as $\Theta$ and $B_{\epsilon}(d_{0})$ are bounded sets, the uniform control over all $\theta,d$ is natural. On the other hand, the linear tilt parameter $t$ is unbounded, so instead of taking a supremum over $t$, we allow the sub-Gaussian norm to depend on $|t|$. This flexibility allows us to cover log-concave as well as non log-concave mixture priors (see Lemma 3.2 below). We note that our results go through with minor modifications if we replace the linear growth of the sub-Gaussian norm in R6 with a general polynomial order growth condition. The linear growth is primarily for expositional simplicity.
The following lemma illustrates concrete instances where conditions R5, R6 in 3.4 hold, which already generalizes existing assumptions in the literature. In particular, we can accommodate priors with unbounded support, mixture priors, and sparsity inducing priors. On the other hand, existing works on low SNR high-dimensional Bayesian regression require the prior to have (a) bounded support [40, 35, 21], (b) log-concave densities [34, 52], or (c) densities that satisfy the log-Sobolev inequality alongside a stronger version of our R4 [22]. These assumptions in the existing literature exclude common choices in practical EB methods, such as adaptive shrinkage priors (this is essentially scale mixtures of Gaussians, see [55]).
@@
    $\mu_{\theta}$ is a finite mixture of $\delta_{0}$ and Lebesgue densities which satisfy (R5), the first condition in (R6), and the stronger second condition in (R6):
@@
    In particular, these conditions for the densities (individual mixture components) are satisfied by distributions from parts (a) and (b) above, so Gaussian mixture priors under any parametrization satisfy R5 and R6.
@@
Part (b) considers higher dimensions with $p\propto n^{2/3}$. While the convergence rate and limiting variance is same as that in part (b), an additional bias appears in the limit. Note that this bias grows as $\delta=\lim_{p\to\infty}\frac{p^{3/2}}{n}$ increases. In part (c), we further increase the dimension so that $\delta\to\infty$, or equivalently $n^{2/3}\ll p$. The resulting limiting distribution exhibits a phase transition and results in a slower rate of convergence (since $\frac{n}{p}\ll\sqrt{p}$ under this regime). This is because now the bias term in part (b) dominates the variance.
To better understand this phase transition, we briefly discuss the proof strategy when $k=1$. Assuming that $\theta$ is a one-dimensional parameter, the first-order conditions for (12) gives
@@
$\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X,\beta^{\star}\approx N\big(\mu_{p}(X,\beta^{\star}),T_{2}\big),~~\text{where}~~\mu_{p}(X,\beta^{\star}):=\frac{1}{\sqrt{p}}\operatorname{\mathbb{E}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X,\beta^{\star}\Big]$ (17)
and $T_{2}>0$ is some constant variance (see (52)). The phase transition on $p$ arises from analyzing this mean term, where we view each summand $F_{i}^{\prime}(\theta_{0})=\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{w_{i},d_{i},\theta_{0}})$ as a function of $w_{i}$ and use second-order Taylor expansion $w_{i}=w_{i}(y)\approx d_{0}\beta_{i}^{\star}$ to write
$\mu_{p}(X,\beta^{\star})\approx\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}\nabla\ell(\theta_{0};B_{d_{0}\beta_{i}^{\star},d_{0},\theta_{0}})+\frac{p^{3/2}}{n}\kappa(\theta_{0})=G_{\theta_{0}}(d_{0}\beta_{i}^{\star})+\frac{p^{3/2}}{n}\kappa(\theta_{0}).$ (18)
Here, the bias term $\kappa(\theta_{0})$ arises from expanding the second term in the Taylor expansion, as the first derivative term turns out to be always negligible.
When $p\ll n^{2/3}$, the first summand in (18) dominates and gives the limit $\mu_{p}(X,\beta^{\star})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,T_{1})$ for another constant $T_{1}>0$ (see (52)). Combining this with the conditional normal limit (17) using the fact that $T_{1}+T_{2}=V(\theta_{0})$, we get $\frac{1}{\sqrt{p}}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}F_{i}^{\prime}(\theta_{0})\mid X\operatorname*{\xrightarrow{\mathnormal{d}}}N(0,V(\theta_{0}))$. The conclusion in part (a) follows by Slutsky’s theorem.
@@
The convergence rate $\sqrt{p}$ is information-theoretically tight. To see this, consider the setting where one directly observes all regression coefficients $\beta^{\star}=(\beta_{1}^{\star},\ldots,\beta_{p}^{\star})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu_{\theta_{0}}$. The resulting convergence rate for estimating $\theta_{0}$ is $\sqrt{p}$, and hence a faster convergence rate is impossible given only $y$.
@@
By applying the Delta method, we can show the following limit for $p=o(n)$:
@@
This limiting variance is sub-optimal and strictly greater than $V(\theta_{0})^{-1}$.
@@
In part (c) of 3.3, note that the $k$-dimensional vector $\kappa(\theta_{0})$ may be zero. While this singularity is typically a measure-zero event (see Figure 1 below for illustration), the convergence rate at such points is faster than the $\frac{n}{p}$-rate in part (c). In particular, if $\kappa(\theta_{0})=0$, we can extend 3.3 to show that $\hat{\theta}^{\mathsf{vEB}}-\theta_{0}=O_{P}\Big(\frac{1}{\sqrt{p}}+\big(\frac{p}{n}\big)^{3/2}\Big)$. Consequently, (16) extends to the higher-dimensional regime of $p=o(n^{3/4})$, and the phase transition no longer happens at $p\asymp n^{2/3}$.
@@
To illustrate the limits in 3.3, we consider a toy example of the Gaussian prior with unknown mean. Here, conjugacy allows directly computing the limit distribution.
@@
Suppose that the prior is $\mu_{\theta}\stackrel{{\scriptstyle d}}{{\equiv}}N(\theta,1)$. Then, the integral for defining the functions $F_{i}$ becomes tractable (see (12)), and we can write
@@
By combining the two displays above and noting $\frac{1_{p}^{\top}\Sigma 1_{p}}{p}=d_{0}+O_{P,X}\Big(\frac{1}{\sqrt{n}}\Big)$ (through standard moment computations), we have obtained the limit distribution of $\hat{\theta}^{\mathsf{vEB}}$: